answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability of your parachute failing?

It depends on how well you pack it, how good a condition it is in, and on how correctly you use it.

In practice, the probability is very low. About one in every 100,000 jumps in the US results in a fatality, a probability of about 0.00001, but that's not just due to parachute failure. One unconfirmed estimate is that one in a thousand (0.001) main chute deployments result in a malfunction, but it is true that very few of those result in a fatality, primarily because there is a reserve chute that is maintained and packed under different, more stringent, requirements.

What is the fifth and 9fifth percentiles of a mean temperature of 98.2 and the standard deviation is 0.62?

The 5th percentile of a standard normal distribution is -1.645 (from the normal probability table).

z = (x - μ) / σ

-1.645 = (x-98.2) / .62

(0.62)(-1.645) = x-98.2

-1.0199 = x-98.2

x = 98.2-1.0199 = 97.1801

The 5th percentile is 97.18

The 95th percentile of a standard normal distribution is 1.645 (from the normal probability table).

z = (x - μ) / σ

1.645 = (x-98.2) / .62

(0.62)(1.645) = x-98.2

1.0199 = x-98.2

x = 98.2+1.0199 = 99.2199

The 95th percentile is 99.22

Does spinning a four section spinner affect a six-sided number cube will land upon when tossed?

Presuming that the spinner and the number cube are both "fair", then no - spinning the spinner and tossing the six-sided number cube are called statistically independent events. They do not influence each other, and it does not matter which order the events occur in.

What is the probibility from 1 out from 27?

If the number is selected at random, then 1/27.

A spinner has the numbers 10 20 30 40 and a number cube numbered 1 through 6 what is the probability of getting a sum greater than 25?

Pr(Sum > 25) = Pr(Spinner = 30 or 40 and Cube = 6)

= Pr(Spinner = 30 or 40)*Pr(Cube = 6)

= 2/4 * 1/6 = 1/2*1/6 = 1/12 or 8.33... %

What would be the chance of picking a multiple of 2 from a deck of cards?

If one card is picked at random from a normal deck of cards, the probability is 20/52 or 5/13.

What was the outcome of this event?

Could you please specify which event you are referring to? This will help me provide a more accurate and relevant answer.

What is sampling error?

Sampling errors are errors in the data collected during the carrying out of quantitative data surveys. They can occur for various reasons, e.g. surveys that were incorrectly filled out. It is generally said that a survey needs to have a margin of error of under 3% to be statistically significant.

What is differential calculus?

That is the part of calculus that is basically concerned about calculating derivatives. A derivative can be understood as the slope of a curve. For example, the line y = 2x has a slope of 2 at any point of the line, while the parabola y = x squared has a slope of 2x at any point of the curve.

What is the difference between double sampling and multiphase sampling?

DOUBLE SAMPLING

Definition:

A standard form of sample design for industrial inspection purposes. In accordance with the characteristics of a particular plan, two samples are drawn, n1 and n2, and the first sample inspected. The batch can then be accepted or rejected upon the results of this inspection or the second sample be inspected and the decision made upon the combined result.

Context:

The term has also been used somewhat loosely for what is called multi-phase sampling and the two-stage version of multi-stage sampling. There is a further usage whereby a first sample provides a preliminary estimate of design parameters which govern the size of the second sample to achieve a desired overall result.

MULTI-PHASE SAMPLING

Definition:

It is sometimes convenient and economical to collect certain items of information from the whole of the units of a sample and other items of usually more detailed information from a sub-sample of the units constituting the original sample. This may be termed two-phase sampling, e.g. if the collection of information concerning variate, y, is relatively expensive, and there exists some other variate, x, correlated with it, which is relatively cheap to investigate, it may be profitable to carry out sampling in two phases.

At the first phase, x is investigated, and the information thus obtained is used either (a) to stratify the population at the second phase, when y is investigated, or (b) as supplementary information at the second phase, a ratio or regression estimate being used.

Two-phase sampling is sometimes called "double sampling".

Context:

Further phases may be added if desired. It may be noted, however, that multiphase sampling does not necessarily imply the use of any relationships between variates x and y. The expression is not to be confused with multi-stage sampling.

What is the probability of winning a lottery game where the winning numbers is made up of seven digits from 0 to 9 chosen at random?

It depends on how the game is structured, a piece of information that was not given in the answer...

If the numbers are reused, then there are 1 x 107 (ten billion, in the short scale) permutations. Without combinations, then, the probability is 1 x 10-7.

If the numbers are not reused, and we are playing combinatorial, in the style of most state sponsored lotteries, then the number of combinations is (10-3)! / 7!, which is 120, making the probability 1 in 120, or about 0.00833.

There are other possibilities but, again, it depends on the rules, and you did not specify them completely.

There are 52 cards in a deck of playing cards and each deck has four sixes If you pick one card from the deck at random what is the probability that you will pick six?

4/52 = 1/13

There are 4 sixes (the numerator)

There are 52 cards (denominator)

Reduce your fraction

You have 4 chances out of 52 that the card you pick will be a six (because there are 4 sixes in the 52-card deck)

What are the features of a binomial distribution?

An experiment with only two outcomes ("success" and "failure"), a constant probability of success, a number of independent trials.

Then, if the probability of a success in a trial is p, the probability of r successes in n trials is

nCr*pr*(1-p)(n-r) for r = 0, 1, 2, ..., n.

In case the super-and sub-scripts do not work, that is

n!/[r!*(n-r)!]*p^r*(1-p)^(n-r)

Explain how to estimate 498 divided by 12?

Let's see. Estimating would start with rounding up 498. 498 will turn into 500. Next, we round 12. Twelve rounds to 10, so it ends up to become 500 divided by 10. 500 divided by 10 is 50. So, estimated, 498 divided by 12 would be somewhere near 50.

(by the way, 498 divided by 12 is actually, non-estimated, 41.5.)

How do you work out the probability that a biased dice will land on a six is 0.35?

You roll it many times. The probability that it lands on a six is the number of times that it lands on a six divided by the number of times the die has been rolled.

What is the least common factor of 60 and 84?

The least common factor of any set of positive integers is 1.

Trending Questions
What are the chances of dying in the Vietnam war? Which distribution is used to find probabilities about the number of independent events occurring in a fixed time period with a known average rate? On the show a thousand ways to die what's the number one way to die? What is the probability of drawing a king of hearts from a regular deck of cards? Strategies you can use to find theoretical probabilities? A womens father is colorblind She marries a colorblind man Will there son or daughter be colorblind? If I were to flip a coin 2 times. What is the probability of getting heads? What if you have a mean standardized score of 1500 points with a standard deviation of 150 points this data is normally distributed What is the z score of 1650? A card with one spot or mark? Can a z score be a negative value? What is the probability of selcting two s's if the first card is not replaced before selcting the second card? What is the difference between the normal distribution and Poisson distribution? How many diagonals does a 13 sided figure have? How many unique ways are there to arrange the letters in the word OPAL? What is an expected outcome in probability? What is the probability of drawing a queen or a club from a deck of cards? When drawing a card from a deck the Pface card and hearts is? What z-score represents the 99th percentile? What does it mean when you look at your watch randomly and the hour and minutes are the same 10.10 12.12 09.09? A researcher polled 230 freshmen at a university and found 110 of them were enrolled in a history class Estimate the probability that a randomly selected freshman is enrolled in a history class?